# TermSight: Making Service Contracts Approachable

### v1

@@ 4.3. Summary Snippets: Plain language summaries of discrete information snippets
To help users surface and interrogate snippets of information buried in a wall of text (DG2) and make original text more accessible (DG4), TermSight provides Summary Snippets accompanied with Highlight Bars for seamless navigation between the summary and original text.
@@ 4.3.1. Summary Snippets
Unlike prior methods that rely on document-level overviews (Tomuro et al., 2016; Keymanesh et al., 2020) or section-level summaries (Manor and Li, 2019; August et al., 2023), TermSight supports interaction and sensemaking for finer-grained snippets of information within dense sections. TermSight features Summary Snippets - one-sentence plain language summaries of Information Snippets (§4.1.1). Each Summary Snippet is rendered with color based on the Power and Relevance of the corresponding Information Snippet (§4.1.2). To enable users to interrogate original text and verify AI-generated summaries (DG4), clicking on a Summary Snippet jumps a user to the corresponding Information Snippet in the document (Figure  4a).
@@ 4.3.2. Highlight Bar
To help users surface and identify Information Snippets while reading or skimming the original text (DG2), TermSight introduces Highlight Bars. Highlight bars are positioned to the left of each Information Snippet in the original text, visually segmenting dense sections into Information Snippets. For paragraphs containing multiple Information Snippets, additional line breaks are added after each snippet to add visual separation. Similar to the Summary Snippets, Highlight Bars are color-coded based on the Power and Relevance of the corresponding Information Snippet (§4.1.2). To support navigation from the original document to the summary, users can click on a Highlight Bar to jump to the corresponding Summary Snippet, as shown in Figure  4b.
Figure 4. Viewport Scrolling: (a) By clicking the Summary Snippet, the viewport of the original passage will scroll to the referenced Information Snippet; (b) By clicking the Highlight bar, the left viewport will scroll to the corresponding Summary Snippet.
Figure 5. Phrase Scope: A tooltip that provides term definition and scenario that contextualizes the meaning of the phrase while allowing users to ask clarification questions. Users can hover over the references to view the referenced text in the ToS.
4.4. Phrase Scope: Context-dependent definition, personalized scenario, and clarification question
To help users read original text with unfamiliar or vague terminology (DG3), TermSight extends prior work focusing on offering dictionary definitions (e.g.,  (August et al., 2023)) by introducing Phrase Scope: a set of features that support the identification and interpretation of unfamiliar or vague phrases (Figure  5). First, TermSight identifies words and phrases that might be unfamiliar or vague to general audience readers and underlines these phrases in blue. This aims to provide a starting point for users to explore potentially unfamiliar phrases without relying on users to identify them. Clicking on the identified phrase opens Phrase Scope, a tooltip with the phrase’s definition in context, a personalized scenario to envision potential implications, and an option to ask clarification questions (examples in Table  2). In addition to the suggested phrases, TermSight allows users to highlight any arbitrary span of text to request Phrase Scope.
@@ 5. Implementation details
TermSight is rendered as a web interface implemented with Next.js. Below, we explain the input, output, processing, and overall performance of TermSight.
5.1. System Input and Pre-processing:
@@
TermSight assumes a set of HTML or markdown source files representing the policies included in the Terms of Service (ToS). For our user study, the research team manually copied the text from documents in ToS on the internet into markdown files with annotated hyperlinks and headers. We assumed clean source files because the focus of TermSight is not on developing web scraping and cleaning technologies but rather on investigating meaningful navigational and reading affordances for ToS. Each document is segmented into text ”chunks”, composed of one or more paragraphs, which are further vectorized and stored in a vector database (detailed in Appendix  B.1).
For the TermSight features that depend on user preferences, such as classifying the relevance of information snippets (§5.2) and generating personalized scenarios (§5.3.3), the system uses a text-based persona that includes users’ intended usage of the service (e.g., content consumers vs. content creators) and their values or concerns (e.g., privacy, copyright, etc.). The user personas used for the study are further explained in Section  6.1.2 and displayed in Appendix  F.1.
Figure 6. A flowchart of the implementation for (1) obtaining Summary Snippets and Information Snippets and (2) classifying Power and Relevance for each Information Snippet.
5.2. Obtaining Summary Snippets and Information Snippets:
@@
For each chunk of text obtained from the document pre-processing pipeline, we prompted GPT-4o to generate a list of 1-sentence summaries (Summary Snippets) each referenced to a span of the input text chunk (Information Snippet) as illustrated in Figure  6. While the length of the referenced text for each summary is unrestricted, each summary is constrained to a maximum of 12 words based on prior works (Cutrell and Guan, 2007; Sweeney and Crestani, 2006) and prompt engineering. Then, we prompted GPT-4o to classify each Information Snippet along two dimensions: Power and Relevance (to user persona). We detail our prompt design in Appendix  B.3.
5.3. Phrase Scope
Below, we detail the implementation of each feature in Phrase Scope.
5.3.1. Identifying unfamiliar or vague phrases
Prior works have demonstrated the capabilities of LLM in personalized jargon identification (Guo et al., 2024) and detecting phrases that can potentially be elaborated (Fok et al., 2024a). Inspired by this work, we employed GPT-4o with two few-shot prompts to identify potentially unfamiliar and vague phrases within the document chunks produced by the pre-processing pipeline (Prompts in Appendix  B.4.1). The input chunks are also referred to and used as the ”phrase context” for each identified phrase.
5.3.2. Generating phrase definitions or responses to user questions
@@
To generate in-context definitions for the identified phrases, we used a retrieval-augmented question answering approach  (Fok et al., 2024a). We first query the vectorized document chunks (§5.1) to retrieve chunks that might define the phrase. The database query is framed as a question: “What does {input phrase} refer to in the sentence: {phrase context}. Here, phrase context refers to the chunk of text containing the input phrase as defined in §5.3.1. Then, the query is embedded using OpenAI’s text-embedding-3-small model, and the top 15 chunks from the vector database are retrieved based on cosine similarity. These chunks are collectively referred to as ”retrieved chunks”. Using the phrase context and the retrieved chunks, we prompted GPT-4o to generate an in-context definition with references to the retrieved chunks. When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked. We detail our prompts in Appendix  B.4.2.
5.3.3. Generating scenarios for phrases
Scenarios intend to complement phrase definitions, helping users interpret and envision the potential implications of a phrase in context. In addition, customizing scenarios based on user personas intends to ensure that the generated examples are relevant to the user’s intended usage of the service (e.g., content creator or consumer) and personal value (e.g., privacy, copyright). Building on prior works that demonstrated the effectiveness of large language models (LLMs) in generating unintended consequences (Wang et al., 2024) and stories to teach legal concepts (Jiang et al., 2024), we employed GPT-4o with zero-shot prompting (Figure  27) to generate the scenarios with the phrase context, phrase definition, and user persona.
@@ 5.4. Evaluation of System Output
Before reporting on our user study, we randomly sampled and conducted a manual evaluation of TermSight’s core outputs. Our goal was not to assess advances in system performance, but to verify that the system can produce meaningful outputs that could support ToS reading. Across 116 sampled information snippets, we found 5 imperfect classifications of power or relevance due to the lack of full context in the input information snippet. Similarly, we reviewed 113 generated definitions and scenarios each. One scenario was found to be factually incorrect due to hallucination. All the definitions were correct except for 4 overly general definitions for service-specific phrases not explicitly defined anywhere in the ToS. More details are provided in Appendix C.
@@ 6. User Study
We conducted a counterbalanced within-subject experiment with 20 participants to evaluate TermSight. We asked the following research questions:
@@
RQ2: How did participants read with TermSight and its features?
RQ3: How did TermSight influence comprehension and recall?
@@ 6.1. Study Design
6.1.1. Study Conditions/Interface Variants
To evaluate TermSight, we conducted a within-subject experiment where each participant used both TermSight and a baseline interface to read 2 services’ ToS contracts, one with each interface variant. The order of the interface variant and service type were counterbalanced to reduce ordering effects. The baseline interface intended to mimic a standard ToS. The baseline interface had the same layout as TermSight without the Power Meter, Summary Snippet, Highlight Bar, and Phrase Scope. In place of the Summary Snippets on the left was a table of contents. Participants could click a section header in the table of contents to navigate to the corresponding section.
6.1.2. Materials: ToS and User Persona
For the user study, we used the ToS of one social media site (Reddit) and one e-commerce site (Poshmark) because social media and e-commerce platforms are among the most visited digital services and are representative of the standard long-form contract used in prior studies  (Kay and Terry, 2010; Taber et al., 2020; Good et al., 2005; Tabassum et al., 2018; Obar and Oeldorf-Hirsch, 2020). For each service, we manually collected policies that are linked or hyperlinked as part of the ToS. We did not include policies that are location-specific (e.g., a California Privacy Notice). For Reddit, we also did not include policies for advertisers, brand&press, or governmental agencies as they were not relevant to a lay user and to keep the number of policies for both services similar. We collected a total of 14 policies for Reddit and 15 for Poshmark. The service names for Reddit and Poshmark were anonymized as ServiceX and ServiceY. Features of TermSight rely on a user persona to determine relevance. For the user study, we designed two personas based on users’ information needs identified in the formative study: a content consumer who posts personal content on social media sites (Figure  17) and a buyer who rarely posts reviews on E-commerce sites (Figure  18).
Figure 7. The experiment procedure for the within-subject study. At the top, we show the procedure for the entire study. At the bottom, we show the process diagram for a reading session with either TermSight or Baseline. The orders for the reading session were counterbalanced in the study.
@@ 6.1.3. Study Procedure
The entire study took 90 minutes and participants were compensated 25\$. The study was approved by the Institutional Review Board (IRB) office at our organization. The study was composed of 2 reading sessions, 45 minutes each (Figure  7). After obtaining consent and before the reading sessions, we asked participants about their past experiences with ToS. Within each reading session, participants first completed a pre-survey and reported their past experiences interacting with social media or e-commerce platforms. Then, an anonymized description of the service and user persona was given. Participants were given 10 minutes to read the ToS without having to speak aloud. Reminders were given at the 5 and 1-minute mark. After the reading session, participants completed seven 5-point Likert-style rating questions about their reading experiences followed by one free-response recall question and six multiple-choice comprehension questions. We also conducted semi-structured interviews with participants about their experiences and reading strategies for 10 minutes. The same procedure was repeated for the second reading session. The study materials are included in Appendix  F and the supplemental materials.
Figure 8. Participants’ self-ratings of their general familiarity with social media and e-commerce platforms and their alignment with the user persona given for each service. Most participants were familiar with both social media and e-commerce platforms in general and found the given persona to align with their personal usage of the service and personal value.
@@ 6.2. Participants and Setting
We recruited 22 participants through Prolific. The recruitment required participants to be over 18, fluent in English, located in the US, not have color deficiencies, and not be legal professionals. We did not require participants to have prior experience in reading ToS. Two participants encountered technical difficulties, and their data were removed from the analysis. Of the remaining 20 participants, 13 self-identify as female and 7 as male. 1 participant had a high school degree, 3 had an associate degree, 12 had a bachelor’s degree, and 4 had a master’s degree. 4 participants were between the ages of 18-25, 5 were between 26-35, 5 were between 36-45, and 6 were between 46-55. When asked about how many ToS they’ve read, 4 participants read none, 5 read between 1—3, 5 read between 4—6, 1 read between 7—9, and 5 read greater than 10. The median of participants’ self-rated familiarity with ToS was 3.5 ($\sigma=1.5$). Most participants were generally familiar with social media and e-commerce platforms and found their given user persona for both services to highly align with their personal usage of the service and their personal value (Figure  8). None of the participants had read or remembered the ToS for Reddit and Poshmark prior to the study.
@@ 6.3. Measures
6.3.1. Perceived reading experiences
We collected 5-point Likert-style ratings (1=“Not at all,” 5=“Very”) of participants’ reading experiences (Appendix  F.3). The experience measures included ease of reading, perceived understanding, and confidence in obtaining relevant information  (August et al., 2023). In addition, we included two 5-point Likert-style rating questions related to the ease of deciding which policy to read and what text to read within a policy, both of which are challenges identified in our formative study. Lastly, we included two 5-point Likert-style questions about their willingness to (1) read other ToS with the interface and (2) spend more time on their given ToS with the interface and get a link to the ToS after the study  (August et al., 2024).
6.3.2. Reading comprehension and recall
For reading comprehension, we wrote 10 multiple-choice questions for each service and selected 6 questions that were relevant for the given user persona for each service, similar to prior works on evaluating reading comprehension for ToS and privacy policy (Tabassum et al., 2018; Vu et al., 2007; August et al., 2023). The questions are different for the two services to minimize learning effects. Comprehension was measured as the number of questions participants got right out of 6. Examples of comprehension questions and their answers are listed in Table  1.
For recall, participants were asked to write down information in ToS they found interesting, surprising, or any detail they remembered: “What did you learn from reading the ToS? Recall one or more interesting things you learned from the ToS.” To analyze these free-form responses, we manually broke the response into single references to a clause in the ToS. Then, we labeled whether each reference was correct based on the ToS. Recall is measured as the number of correct facts participants wrote in the response  (Taber et al., 2020; Good et al., 2005).

### v2

@@ 4.3. Summary Snippets: Plain language Summaries of Information Snippets
To help users surface relevant information within a document (DG2) and interrogate the original text (DG4), TermSight provides Summary Snippets accompanied by Highlight Bars for seamless navigation between the summary and the original text (Figure  1).
@@ 4.3.1. Summary Snippets
In contrast with document-level  (Tomuro et al., 2016; Keymanesh et al., 2020) or section-level summaries (Manor and Li, 2019; August et al., 2023), TermSight explores supporting interaction and sensemaking for finer-grained snippets of information within dense sections. TermSight features Summary Snippets: one-sentence plain language summaries of Information Snippets (§4.1.1). Summary Snippets are designed to be short (detailed in Section 5.2), which have been shown to be more effective for navigational tasks compared to longer snippets  (Cutrell and Guan, 2007; Sweeney and Crestani, 2006). Each Summary Snippet is rendered with color based on the Power and Relevance of the corresponding Information Snippet to provide information scent (Pirolli and Card, 1999) for deciding which snippets to pay attention to (DG2). To enable users to interrogate original text and verify AI-generated summaries (DG4), clicking on a Summary Snippet jumps a user to the corresponding Information Snippet in the document.
@@ 4.3.2. Highlight Bar
To help users surface and identify Information Snippets while reading or skimming the original text (DG2), TermSight introduces Highlight Bars. Highlight bars are positioned to the left of each Information Snippet in the original text, visually segmenting dense sections into Information Snippets. For paragraphs containing multiple Information Snippets, additional line breaks are added after each snippet for visual separation. Similar to the Summary Snippets, Highlight Bars are color-coded based on the Power and Relevance of the corresponding Information Snippet (§4.1.2). Users can click on a Highlight Bar to jump to the corresponding Summary Snippet.
Figure 4. Phrase Scope first identifies jargon or vague phrases in the contract (a). Clicking an identified phrase opens a tooltip with definition (b) and hypothetical scenario (c) while allowing users to ask clarification questions (d).
Phrase Scope, a tooltip that provides a term definition and a scenario to contextualize the meaning of a phrase, while also allowing users to ask clarification questions. Users can access Phrase Scope by clicking on a system-identified phrase or highlighting any phrase they do not understand. The figure shows an example output for the phrase ‘royalty-free’. Definition: ‘Royalty-free refers to the license to use, copy, modify, or distribute your content without requiring licensing fees.’ Scenario: ‘Imagine you post a photo on ServiceX, and they use it in a global ad campaign without paying you. Because you agreed to a royalty-free license, ServiceX can freely use your content for commercial purposes without compensating you, reducing your control and potential earnings from your creation.’
4.4. Phrase Scope: Phrase Identification, Definition, Scenario, and Clarification
To help users engage with the original text (DG3), Phrase Scope first identifies phrases that may be unfamiliar or ambiguous to general audience readers and underlines them in blue (Figure 4). This provides a starting point for users to explore potentially unfamiliar phrases without relying on users to identify them. Users can also highlight any arbitrary span of text to request Phrase Scope. Clicking on the system or user-identified phrase opens Phrase Scope, a tooltip with three components. First, the definition of the phrase in the context of the ToS is offered, similar to prior work in other domains (Fok et al., 2024a). Beyond definitions, participants in the formative study described the challenge to contextualize abstract phrases in unforeseen scenarios. Inspired by prior work that encouraged scientists to reflect on the unintended consequences of their research  (Wang et al., 2024; Pang et al., 2024), Phrase Scope explores the opportunity to prompt reflection on the implications of a phrase by offering hypothetical scenarios personalized to the user persona. Lastly, users have an option to ask clarification questions. More examples of the generated definitions and scenarios can be found in Appendix Table  2.
@@ 5. Implementation Details
TermSight is rendered as a web interface implemented with Next.js. Below, we explain the input, output, processing, and overall performance of TermSight. Additional implementation details and prompts are specified in Appendix  B.
5.1. System Input and Pre-processing
@@
TermSight assumes a set of HTML or markdown source files representing the policies included in the Terms of Service (ToS). We assumed clean source files because the focus of TermSight is not on developing web scraping technologies but rather on investigating meaningful navigational and reading affordances for ToS contracts. Each document is segmented into text ”chunks”, composed of one or more paragraphs, which are further vectorized and stored in a vector database (detailed in Appendix  B.1). For features that depend on user preferences, such as classifying the relevance of Information Snippets (§5.2) and generating personalized scenarios (§5.3), TermSight uses a text-based persona that includes users’ intended usage of the service (e.g., content consumers vs. content creators) and their values or concerns (e.g., privacy, copyright, etc). The user personas used for the study are further explained in §6.1.2 and displayed in Appendix  E.2.
5.2. Obtaining Summary Snippets and Classifying Information Snippets
@@
For each chunk of text obtained from the document pre-processing pipeline, we prompted GPT-4o to generate a list of one-sentence summaries (Summary Snippets) each referenced to a span of the input text chunk (Information Snippet) as illustrated in Figure  5. While the length of the referenced text for each summary is unrestricted, each summary is constrained to a maximum of 12 words based on prompt engineering and prior work showing that short summaries can be more effective for navigation (e.g., 10-20 words)  (Cutrell and Guan, 2007; Sweeney and Crestani, 2006). We noticed that setting summaries shorter than 12 words led to excessive fragmentation (i.e., too many Summary Snippets), while longer summaries tried to cover too much information, making them less skimmable. We then prompted GPT-4o to classify each Information Snippet along two dimensions: Power and Relevance (to user persona). Prompts can be found in Appendix  B.2 and  B.3.
Figure 5. A flowchart of the implementation for (1) obtaining a list of Summary Snippets each referenced to a span of the input text (i.e., Information Snippet) and (2) classifying Power and Relevance for each Information Snippet.
The flowchart of TermSight’s implementation process. (1) The figure shows that a chunk of text is first summarized into a list of Summary Snippets, each referenced to a span of the input text chunk (Information Snippet). (2) After this step, TermSight classifies every Information Snippet with respect to Power and Relevance.
5.3. Generating Phrase Scope
@@
Similar to prior work (Guo et al., 2024; Fok et al., 2024a), we prompt an LM to identify potentially unfamiliar and ambiguous phrases within document chunks produced by the pre-processing pipeline. To generate in-context definitions for the identified phrases, we used a retrieval-augmented question answering approach  (Fok et al., 2024a). We vectorize the document chunks (§5.1) and query these chunks to retrieve relevant document context. These chunks are collectively referred to as ”retrieved chunks”. Then, we prompted GPT-4o to generate an in-context definition with references to the retrieved chunks. When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked. Scenarios aim to prompt reflection on the potential implications of a phrase based on users’ intended usage of the service and personal value. We employed GPT-4o with zero-shot prompting to generate the scenarios tailored to the user persona. The prompts used for Phrase Scope are detailed in Appendix  B.4.
@@ 5.4. Evaluation of System Output
Before reporting on our user study, we randomly sampled and conducted a manual evaluation of TermSight’s core outputs (e.g., summary snippets, definitions, scenarios). Our goal was not to assess advances in system performance, but to verify that the system can produce meaningful outputs that could support ToS reading. Across 116 sampled Information Snippets, we found 5 imperfect classifications of power or relevance due to the lack of full context in the input Information Snippet. Similarly, we reviewed 113 generated definitions and scenarios each. One scenario was found to be factually incorrect due to a hallucination. The input context and phrase stated that users do not gain ownership rights by downloading content from the service. However, the scenario claimed that users might lose ownership rights over the content they create by uploading it to the service. All the definitions were correct except for 4 overly general definitions for service-specific phrases not explicitly defined anywhere in the ToS. More details of the evaluation are provided in Appendix C.
@@ 6. User Study
Similar to prior HCI studies  (Chen et al., 2023; Cao et al., 2023; Kambhamettu et al., 2024), the goal of our user study is not to prove TermSight as the definitive solution through a series of ablation experiments, but to use TermSight to explore the problem and solution space of designing contract reading interfaces. We asked the following research questions:
@@
RQ2: How did TermSight influence ToS comprehension and recall?
RQ3: How did participants read with TermSight and its features?
@@ 6.1. Study Design
We conducted a within-subject study where each participant used both TermSight and a baseline interface to read two services’ ToS, once with each interface variant. The order of the interface and service type was counterbalanced to reduce ordering effects.
6.1.1. Baseline
Because there is no well-established standard ToS reader, we use the most commonly used interface as the baseline (i.e., an HTML reader). The baseline interface had the same layout as TermSight without the Power Meter, Summary Snippet, Highlight Bar, and Phrase Scope (Appendix Figure  25). In place of the Summary Snippets was a table of contents that can be used to navigate to different sections. Participants’ feedback in the user study confirmed that the baseline offered a strong frame of reference (§7.1). While reading interfaces exist for other domains (e.g.,  (August et al., 2023; Chen et al., 2023; Fok et al., 2023)), adopting them to contracts would result in novel systems unsuitable as baselines. For example, prior work has designed key questions (August et al., 2023) or faceted highlighting (Fok et al., 2023) specific to academic papers, which require co-design studies with lawyers to adapt them to contracts. Injecting additional features (e.g., overview summaries  (Tabassum et al., 2018), chat (Zhao and Lee, 2020)) into one or both conditions may also introduce confounds unrelated to our research questions. For example, by adding a policy or section-level summary in the baseline, the comparison with TermSight will be confounded by the design of summaries in the baseline (e.g., levels of complexity and detail), a variable that may impact different readers differently (August et al., 2024).
6.1.2. Materials
Terms of Service
For the user study, we used the ToS of one social media site (Reddit) and one e-commerce site (Poshmark) because social media and e-commerce platforms are among the most visited digital services and are representative of the standard long-form contract used in prior studies  (Kay and Terry, 2010; Taber et al., 2020; Good et al., 2005; Tabassum et al., 2018; Obar and Oeldorf-Hirsch, 2020). For each service, we collected policies that are linked as part of the ToS. We did not include location-specific policies (e.g., a California Privacy Notice). For Reddit, we also did not include policies for advertisers, publishers, or governmental agencies to keep the number of policies for both services similar. We collected a total of 14 policies for Reddit and 15 for Poshmark. The service names for Reddit and Poshmark were anonymized as ServiceX and ServiceY.
User Persona
Features of TermSight rely on a user persona to determine relevance. For the user study, we designed two personas based on users’ information needs identified in the formative study: a content consumer who posts personal content on social media sites and a buyer who rarely posts reviews on e-commerce sites (Appendix  E.2). Participants in the user study validated and found their given user persona for both services to highly align with their personal usage of the service and their personal value (Figure  13).
@@ 6.1.3. Study Procedure
The entire study took 90 minutes, and participants were compensated 25\$. The study was approved by the Institutional Review Board (IRB) at our organization. The study was composed of two reading sessions, 45 minutes each. After obtaining consent and before the reading sessions, we asked participants about their past experiences with ToS. Within each reading session, participants first completed a pre-survey and reported their past experiences interacting with social media or e-commerce platforms. Then, an anonymized description of the service and user persona was given. Similar to prior studies (August et al., 2023), participants were given 10 minutes to read the ToS without having to speak aloud. Reminders were given at the 5 and 1-minute mark. After the reading session, participants completed seven 5-point Likert-style rating questions about their reading experiences, followed by one free-response recall question and six multiple-choice comprehension questions. We also conducted semi-structured interviews with participants about their experiences and reading strategies for 10 minutes. The same procedure was repeated for the second reading session. The study materials are included in Appendix  E.
@@ 6.2. Participants and Setting
We recruited 22 participants through Prolific. The recruitment required participants to be over 18, fluent in English, located in the US, not have color deficiencies, and not be legal professionals. We did not require participants to have prior experience in reading ToS. Two participants encountered technical difficulties, and their data were removed from the analysis. The median of participants’ self-rated familiarity with ToS was 3.5 ($\sigma=1.5$) out of 5. Most participants were generally familiar with social media and e-commerce platforms and found their given user persona to highly align with their own profile (Figure  13). None of the participants had read or remembered the ToS for Reddit and Poshmark prior to the study. Additional participant demographics can be found in Appendix  E.3.
@@ 6.3. Measures
6.3.1. Reading Experiences
We collected seven 5-point Likert-style ratings (1=“Not at all,” 5=“Very”) of participants’ reading experiences (Appendix  E.5). The experience measures included ease of reading (adopted from perceived effort measure in NASA TLX)  (August et al., 2023; Hart and Staveland, 1988), perceived understanding  (August et al., 2023), and confidence in obtaining relevant information  (August et al., 2023). We added two questions specific to challenges identified in our formative study: the ease of deciding which policy to read and what text to read within a policy. Finally, we included two questions on participants’ willingness to read ToS with the interface and their willingness to spend more time on the ToS with the interface (August et al., 2024).
6.3.2. Comprehension
We wrote 10 multiple-choice questions for each service and selected 6 questions that are relevant for the given user persona (§6.1.2) for each service, similar to prior work on evaluating reading comprehension for ToS and privacy policies (Tabassum et al., 2018; Vu et al., 2007; August et al., 2023). We designed the comprehension questions to have one answer based on the original text, independent of features of TermSight (e.g., Summary Snippets). Consequently, participants need to read the original text to fully answer the questions. The questions were different for the two services to minimize learning effects. Comprehension was measured as the number of questions participants got right. We note that past studies with reading interfaces often fail to show differences in how people answer comprehension questions (Tabassum et al., 2018; Korunovska et al., 2020; Robinson and Zhu, 2020; August et al., 2023; Head et al., 2021). Similar to prior work  (August et al., 2023), our intention in including comprehension questions was to ensure that TermSight did not detract from overall comprehension. The questions can be found in the supplemental materials.
6.3.3. Recall
For recall, participants were asked to write down information in ToS they found interesting, surprising, or any detail they remembered: “What did you learn from reading the ToS? Recall one or more interesting things you learned from the ToS.” To analyze these free-form responses, we manually broke the response into single references to a clause in the ToS. Then, we labeled whether each reference was correct based on the ToS. Recall is measured as the number of correct facts in the response  (Taber et al., 2020; Good et al., 2005).